OpenStudy (raegan360brown):

cos(theta) sin(theta)+2cos(theta)=0

jimthompson5910 (jim_thompson5910):

\[\Large \cos(\theta)\sin(\theta)+2\cos(\theta)=0\] \[\Large \cos(\theta)\left[\sin(\theta)+2\right]=0\] \[\Large \cos(\theta) = 0 \ \text{ or } \ \sin(\theta)+2=0\] Do you see how to finish up?

OpenStudy (raegan360brown):

i typed this problem in wrong there is suppose to be a 7 in front of the first cosine

OpenStudy (raegan360brown):

7cos(theta) sin(theta)+2cos(theta)=0 thats the problem

OpenStudy (triciaal):

sin range is from -1 to 1

jimthompson5910 (jim_thompson5910):

\[\Large 7\cos(\theta)\sin(\theta)+2\cos(\theta)=0\] \[\Large \cos(\theta)\left[7\sin(\theta)+2\right]=0\] \[\Large \cos(\theta) = 0 \ \text{ or } \ 7\sin(\theta)+2=0\]

OpenStudy (raegan360brown):

but when i typed it in my homework it said that -2/7, which it ask for the decimal, is -.286 and it says its wrong

jimthompson5910 (jim_thompson5910):

-2/7 isn't the answer

OpenStudy (raegan360brown):

i got the answers for the cos right though but i dont understand how 3.431+2kpi and 5.993+2kpi is the answer

jimthompson5910 (jim_thompson5910):

\[\Large 7\sin(\theta)+2=0\] \[\Large 7\sin(\theta)=-2\] \[\Large \sin(\theta)=-\frac{2}{7}\] Now use arcsine to isolate theta

OpenStudy (raegan360brown):

so theta = -.290?

jimthompson5910 (jim_thompson5910):

add on 2pi to that (2pi is roughly 6.28)

jimthompson5910 (jim_thompson5910):

arcsin(-2/7) = -0.28975170143604 -0.28975170143604 + 2pi = 5.99343360574354 so that accounts for one of the answers

OpenStudy (raegan360brown):

wait so im supposed to add 2pi even though at the end you add 2pik ?

jimthompson5910 (jim_thompson5910):

the 2pi*k takes care of all coterminal angles (k is any integer)

jimthompson5910 (jim_thompson5910):

`-0.28975170143604 radians` and ` 5.99343360574354 radians` are coterminal angles

OpenStudy (raegan360brown):

oh i understand but how do you know to add the 2pi ?

jimthompson5910 (jim_thompson5910):

because to find a coterminal angle, you add or subtract 2pi because there are 2pi radians in a full circle

OpenStudy (raegan360brown):

oh okay i understand now! thank you so much!

jimthompson5910 (jim_thompson5910):

no problem